OpenStudy (anonymous):

What is the GCF of fifteen and 9? 1 3 5 9

OpenStudy (anonymous):

Do you know what GCF stands for?

OpenStudy (anonymous):

I know I know!

OpenStudy (anonymous):

lol not you @OrangeMaster

OpenStudy (anonymous):

Greatest Common Factor

OpenStudy (anonymous):

I knew that ^

OpenStudy (anonymous):

lol, you should hon

OpenStudy (anonymous):

Yes. I know

OpenStudy (anonymous):

What is the greatest number that you can divide both 9 and 15 by.

OpenStudy (anonymous):

What factor do both numbers share?

OpenStudy (anonymous):

Exactly ^

OpenStudy (anonymous):

9 cannot go into 15 - rule that out Nor is is 1

OpenStudy (anonymous):

5 cannot go into 9

OpenStudy (anonymous):

But 3x3 and 3x5 are legible

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Do you know the answer. @kryssiibaee

OpenStudy (anonymous):

So, your answer is.............

OpenStudy (anonymous):

4

OpenStudy (anonymous):

3 ;)

OpenStudy (anonymous):

4 isn't an option, neither can it go into 15.

OpenStudy (anonymous):

So yeah, the answer is 3.

OpenStudy (anonymous):

wait i was looking at the wrong one. hold on .

OpenStudy (anonymous):

Don't guess.

OpenStudy (anonymous):

Do you have the answer @kryssiibaee and know why it's the correct answer.

OpenStudy (anonymous):

Yes. But can you explain something else to me ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

okay can you explain why this isnt 4\16

OpenStudy (anonymous):

What do you mean? ^

OpenStudy (anonymous):

i mean 2\8

OpenStudy (anonymous):

What is four-sixteenths in simplest form? one-eighth one-fourth one-half two-eighths

OpenStudy (anonymous):

That's because 2/8 can be simplified further.

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

Do you get it now, if so what's the right answer.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

1\4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

4/16 Both are divisible by 2 2/8 Again

OpenStudy (anonymous):

Correct honey! :D

OpenStudy (anonymous):

2/8 divisible by 2 1/4! :D

OpenStudy (anonymous):

One more i need you to explain..

OpenStudy (anonymous):

OK, what is it.

OpenStudy (anonymous):

Ok, yeah. What is it?

OpenStudy (anonymous):

I gtg. Hope you get your help @kryssiibaee and a medal for you @OrangeMaster for being so dang awesome....Bai lovelies!!! :D

OpenStudy (anonymous):

Thanks, you're awesome too. ^

OpenStudy (anonymous):

Drew notices there are twenty-one girls in a class of thirty students. What is the simplest form of twenty-one thirtieths? three-tenths three-fifths three-sevenths seven-tenths

OpenStudy (anonymous):

this is 3\7 ?

OpenStudy (anonymous):

orange ?

OpenStudy (anonymous):

No, it would be seven-tenths.

OpenStudy (anonymous):

21/3 and 30/3

OpenStudy (anonymous):

Is that all that you needed?

OpenStudy (anonymous):

are you sure it can't get simplified smaller?

OpenStudy (anonymous):

7 -- can not get smaller, because 7 cannot be divided by anything else 10

OpenStudy (anonymous):

21 / 30 is divisible by 3 7/10 7 is a prime number, so it can only be divided by one So the correct answer is seven tenths or 7/10 like @OrangeMaster said

OpenStudy (zzr0ck3r):

15=9*1+6 9=6*1+3 6=3*2+0 so the gcd(9,15) = 3

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Its GCF and we already got that hun

OpenStudy (zzr0ck3r):

you can use this algorithm to find any gcf

OpenStudy (anonymous):

We see.......

OpenStudy (zzr0ck3r):

gcf = gcd divisor = factor

OpenStudy (anonymous):

If you ever simplify something and end up with prime number, then you know that it is already at the simplest form. Prime Number: A number whose factors are 1 and it's self.

OpenStudy (anonymous):

Which fraction is equivalent to three over ten? fifteen over fifty fifteen over twenty twelve over fifty twelve over twenty

OpenStudy (anonymous):

A

OpenStudy (anonymous):

5*3 = 15 5 * 10 = 50

OpenStudy (zzr0ck3r):

you should just send the hw to them in an email, then they can do it and fax it back:)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

That's hilarious @zzr0ck3r

OpenStudy (anonymous):

Sorry rephrase please. I am sleepy cannot understand

OpenStudy (anonymous):

FAX???? I only have a printer

OpenStudy (anonymous):

zz is saying that we are doing their homework for them.

OpenStudy (anonymous):

He just said it in a hilarious way.

OpenStudy (anonymous):

No, we are aiding. Did we ever give a direct answer? No. We broke it down. Don't be mad you are late and @OrangeMaster is ahead

OpenStudy (anonymous):

Please no hate @zzr0ck3r

OpenStudy (anonymous):

You got it hon?

OpenStudy (anonymous):

Yeah, we did not give any direct answers.

OpenStudy (anonymous):

@kryssiibaee

OpenStudy (anonymous):

Can someone please explain this other question?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Since 5*3 is 15 And 5*10 is 50

OpenStudy (anonymous):

They must use the same numebrs

OpenStudy (anonymous):

lol, this is so cute. I have exactly 30 medals in Math hehe

OpenStudy (anonymous):

Which fraction is equivalent to three over ten? fifteen over fifty fifteen over twenty twelve over fifty twelve over twenty

OpenStudy (zzr0ck3r):

Which fraction is equivalent to three over ten? fifteen over fifty fifteen over twenty twelve over fifty twelve over twenty a)... b)... c).... d).... ... next post Saphira_Shapiro A

OpenStudy (anonymous):

A - Since both are multiplied by 5

OpenStudy (zzr0ck3r):

im not hating, I just saw that and made a joke

OpenStudy (zzr0ck3r):

ignore me:)

OpenStudy (anonymous):

Fifteen over Fifty

OpenStudy (anonymous):

Oh, ok honey @zzr0ck3r

OpenStudy (anonymous):

@kryssiibaee, is that all of the questions?

OpenStudy (anonymous):

I was confused. Thought the sleep was getting to me when I saw the word fax O.o

OpenStudy (anonymous):

I hope so tired by desperately want to help.

OpenStudy (anonymous):

-Tired-But want to help

OpenStudy (anonymous):

@kryssiibaee Any more questions?

OpenStudy (anonymous):

yeah but idk if i should bother you or not.,.

OpenStudy (anonymous):

It's fine, how any more?

OpenStudy (anonymous):

Oh shoot, I wish you asked yesterday when I was up til 2 am hehe

OpenStudy (anonymous):

It's fine, how any more?